OpenStudy (anonymous):

Find the exact value of the following: sin^2(75º)/(sin^45cos^230º)

OpenStudy (anonymous):

\[0.125*(1+\sqrt{3})^2 \csc^{45}(\cos^{230} (\pi/180))\] which is approximately 6191

OpenStudy (anonymous):

thanks so much!

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

sorry, i think i may have written the question unclearly. it's: sin^2(75º)/(sin^45cos^2(30º))

OpenStudy (anonymous):

*Sighs* in that case:\[1/8 (1 + \sqrt{3})^2 \csc^{45}(3/4)\] which is approximately: \[2.882 * 10 ^7\]

OpenStudy (anonymous):

lol sorry, and thanks again.

OpenStudy (anonymous):

it doesn't matter, I'm here to help people :D